# Strategic Planning and Rationalizing on Trees Make LLMs Better Debaters

### v1

@@ 4.1 Experiment and Evaluation Setup
We use the state-of-the-art multi-agent framework Agent4Debate (Zhang et al., 2024) with different backbone LLMs as our baseline. It employs a collaborative architecture with four specialized agents: searcher, analyzer, writer, and reviewer agents. It shows better performance in the LLM-based Debatrix debate metric (Liang et al., 2024) and is preferred by human experts in pairwise comparison. It uses the stage-specific prompts from the AIDebater 2024 competition  [^2] http://www.fudan-disc.com/sharedtask/AIDebater24 and uses Tavily  [^3] https://tavily.com as a search engine for evidence retrieval.
@@
We use Gemini-2.0-flash (Team et al., 2023) and DeepSeek-V3 (Liu et al., 2024) as the backbone LLM for Agent4Debate and TreeDebater. For a fair comparison, TreeDebater uses the same Tavily APIs for evidence retrieval before the debate and adopts the same stage-specific prompts, adding the necessary modifications to incorporate the tree information for planning, as shown in the Appendix G. To ensure that the debate systems follow the time constraint, we add the rough word budget in the stage-specific prompt for both models, which is 520 words for the opening and rebuttal and 260 words for the closing stage. After transferring the statement to the audio via OpenAI TTS, we apply a hard cut on the debate audio: if the audio transferred from the statement exceeds the stage time, it will be trimmed from the last sentence before the time limitation. The statement will also be trimmed accordingly. The other implementation details, such as the attack and defense score functions, are in the Appendix B.3.
@@
Head-to-Head Human Evaluation for Stage-level Comparison. In competitive debates, winning is affected by many factors, such as motions, stance assignments, and the opponent’s performance. This introduces plenty of noise into the evaluation. Therefore, we design a controlled head-to-head setting for stage-level comparison. We first have a debate competition between the two debate systems with a random stance assignment. Then we use the debate process before a target stage as the context, and let each debate system generate one version for this target stage. We then let the participant compare the two versions based on the debate process. For example, we keep the same Pro’s Opening and ask Agent4Debate and TreeDebater to generate the Con’s Opening. This head-to-head comparison provides a detailed and focused evaluation of each stage by alleviating the noise introduced by other factors. Specifically, the annotator will provide persuasiveness scores for both versions and also mark which side performs better in each stage.
@@
End-to-End Human Evaluation for Debate-level Comparison. In addition, we follow Oxford-style debating to assess the debater’s capability in a full debate. We ask the participants to vote before and after listening to the full 20-minute debate, and provide the persuasiveness score for each stage. We flip the stance assignment and take the average of two competitions for each motion to avoid the potential prior bias on the motion.
@@
Before human evaluation, we perform several validity checks to assess the quality of the debate competition generated. Only those that follow the correct debate format are used for human evaluation. The results of the valid checks are shown in Appendix E. In general, TreeDebater can always generate format-valid and time-valid statements. However, only 77% debates are valid in the Gemini version of Agent4Debate. Besides, Agent4Debate always exceeds the time limitation, especially in the closing stage. We put the motion list (Table 4), persuasiveness score rubric (Figure 5), the recruitment details, and a screenshot of our debate evaluation platform (Figure 6) in the Appendix C.
@@
Table 1: Scalar persuasiveness score and win rate in head-to-head human evaluation. A higher score indicates the statement is more persuasive. Win:Tie:Lose indicates the number of cases that TreeDebater wins, gets a tie, or loses in the pairwise comparison. The standard deviation results are put in the appendix.
@@
4.2 Head-to-Head Human Evaluation Results for Single Stage
We randomly sample 10 (motion, stance assignment) settings for each stage and get the corresponding debate context. Then we let each debate system generate one statement for the target stage. The audio of the debate context and the two versions of the statement are presented to the participants, where we randomly set the order of these two versions to avoid the potential order bias. The results are shown in Table 1.
TreeDebater has a higher average persuasiveness and win rate in 11/12 stage-level comparison. The advantage is more significant when we shift from Gemini to DeepSeek. We can find a clearer win in the stage-level win rate. When participants are asked to choose the better version, TreeDebater is preferred 1.5x and 2.5x times than the baseline. One reason could be that participants are more likely to choose scores of 3 or 4, leading to a small performance gap in the persuasiveness score.
4.3 End-to-End Human Evaluation Results for the Full Debate
We randomly sample 4 motions for Gemini and 7 motions for DeepSeek. We conduct two debates on each motion with the original and swapped sides. For each debate, 3 participants are asked to score each stage and vote before and after the debate. The average score and the opinion shift are used to evaluate the debaters’ performance in this debate. We take the average of the two flipped stance assignments as the debater’s performance on this motion. For example, the opening persuasive score is the average of the opening scores when the debaters acted as the Pro and Con side, respectively. Results are shown in Table 2.
TreeDebater outperforms the baseline with a 3.5x and 1.3x higher opinion shift win rate in Gemini and DeepSeek. This indicates that TreeDebater makes the LLM a more persuasive debater that can change the audience’s opinion in the competitive debate. Besides, the debate systems with the DeepSeek backbone have higher persuasiveness scores than Gemini, indicating that DeepSeek has superior capability in argumentation. Moreover, our TreeDebater shows a significant improvement in the persuasiveness score at all stages. We find that this is because the poor performance of the Gemini-based baseline in the early stage affects the participants’ impression of it and makes them more critical in the following stage. On the other hand, we find that the audience focuses more on the personal belief in the motion rather than the debate strategies if the two sides have reasonably good performance in the debate. For example, the Agent4Debate baseline already shows an average persuasiveness score of 3 or more. In its pairwise comparison with TreeDebater, we find that in the 7 motions we used in our DeepSeek experiments, the Con side always wins in 3 of them and the Pro side always wins in one of them, regardless of the stance assignment. This leads to a comparable persuasiveness score after taking the average of the two flipped assignments, as shown in Table 2. In such a scenario, the head-to-head evaluation provides more insight into the performance difference by focusing more on the debate strategies used under the same debate context.
@@ 4.3 End-to-End Human Evaluation Results for the Full Debate
Gemini Agent4Debate 2.96 2.92 2.98 0.13
TreeDebater 3.60 3.58 3.54 0.46
DeepSeek Agent4Debate 3.73 3.81 3.57 0.30
TreeDebater 3.87 3.71 3.38 0.40
@@ 4.4 Analysis
[…]
Figure 2: Action types distribution in the rebuttal stage. We extract the debate flow tree from human debates and categorize the action distribution. The actions are less diverse in the baseline and TreeDebater w/o $T_{d}$.
[…]
Figure 3: Percentage of actions that can be found in the rehearsal trees. The rehearsal trees of both sides contribute to the hit rate.
@@ 5 Conclusion and Discussion
Broader Impacts and Limitation This paper introduces TreeDebater to enhance LLM persuasiveness in competitive debates, with broader potential to improve information efficiency and strategic planning capabilities in games lacking clear winning conditions. However, the study’s conclusions are potentially limited by the scale of human evaluations and the inherent subjectivity of debate performance assessment, despite the authors’ best efforts to mitigate bias through comparative setups. Future work could involve larger, more diverse evaluation groups.
@@ Acknowledgments
This work was supported in part by NSF grant IIS-2046640 (CAREER). This project is also supported by the OpenAI gift and the Gemini Academic Program. Xinran Zhao is funded by the ONR Award N000142312840. The authors thank the Mandarin debate team at HKUST for providing feedback on the debate system and human expert debaters, including Kexun Zhang, Wenting Yang, Yujia Huang, Yue Zhang, Yanran Bao, and Chinmaya Leli, for their valuable feedback.
@@ Appendix A Debate Terminology
We introduce the debate terminologies used in this paper. A motion in the debate is a statement of opinion or proposition to argue. The assigned stance $s$ on the debate motion can be Pro or Con, indicating supporting or opposing the motion, respectively. A claim $c$ is a proposition or assertion about a motion, which should be proved through argumentation. The main claims refer to the most important claims that the debater proposes during the opening stage to support their stance. An argument $x$ is a set of reasons or evidence to support a claim, typically consisting of premises that lead to a conclusion. A statement is the whole passage presented in one debate stage, including a set of claims, arguments, and supporting evidence  [^4] It is also called a speech (Slonim et al., 2021) or a discourse (Zhang et al., 2016) in some literature.. There are several typical actions that the debater can take in each stage: propose a new claim, reinforce its existing claims, attack the opponent’s claims, and rebut the opponent’s attack. A battlefield refers to a conflict zone in the debate, where two sides contest on a specific point. It includes several rounds of propose attack defense.
@@ B.1 Debate Flow Tree Algorithms
2if $action$ is "propose" then
3       Create node $x^{\prime}$ with $claim$ and $argu$;
@@ 5       Modify $x^{\prime}$ status to "proposed" ;
8Find node $x$ similar to $target$ by $\theta$ ;
9 Increase $x$ visit count by $1$ ;
@@
11if $action$ is "reinforce" then
12       Update $x$ with $argu$ ;
@@
15if $action$ is "rebut" or "attack" then
16       Create node $x^{\prime}$ with $claim$ and $argu$;
@@ 19       Modify $x^{\prime}$ status to "proposed" ;
Algorithm 2 Update Debate Flow Tree
@@ 5       Initialize retrieval set $R$ ;
7             if $action$ is "propose" or "reinforce" then
8                   Find node $x^{l}$ similar to $target$ on the current side by $\theta$;
@@ 9                   Add $x^{l}$ and its $k$-step lookahead strength $f_{k}(x^{l})$ to $R$ if exists ;
11             else if $action$ is "attack" or "rebut" then
12                   Find node $x^{l}$ similar to $target$ on the opposite side by $\theta$;
@@ 14                         Add $x^{l+1}$ and its $k$-step lookahead strength $f_{k}(x^{l+1})$ to $R$ if exists ;
16      Add the tuple $action$, $target$, $R$ to $\mathcal{A}_{R}$
@@ B.2 Search Algorithm in Speech Time Control
As mentioned in the main text, our goal is to generate a coherent speech within the time range $[t_{l},t_{r}]$ with the chosen actions from the debate flow tree. Note that the number of words is easier to control than the entire speech length for the large language model generation, and the passage word length has a relatively good correlation with the speech length. We can formulate the problem as that we want to call $f(n)$ minimal times to find an output that has speech time in the range $[t_{l},t_{r}]$, where $f$ represents an LLM inference where the speech length of $f(n)$ is correlated to $n$.
Note that if at any time the generated speech time ranges in $[t_{l},t_{r}]$, we can directly finish the whole process. In our implementation, we first try to establish the boundary word counts for binary search $[n_{l},n_{r}]$, where $n_{l}$ produces a speech shorter than $t_{l}$ and $n_{r}$ produces one longer than $t_{r}$. Specifically, we make an initial query that can either become $l$ or $r$, so that we can find the other endpoint by iteratively doubling $r-l$. With this initial interval determined, we can then iteratively refine this interval through binary search, selecting the midpoint and reducing the interval range by half based on the resulting speech duration.
@@ B.3 Reward Model Training Details
We train two separate LLaMA-based reward models, specifically with the base model being meta-llama/Llama-3.2-3B-Instruct, using the Kialo dataset (Durmus et al., 2019) for $r_{s}$ and $r_{a}$. The Kialo dataset consists of claim texts for 741 controversial motions. For each motion, it maintains an argument tree. The root of the tree is the motion, and each node is a claim text with its counter-arguments or subsequent arguments as its children. There are human annotations for the impact score (Impactful, Medium Impactful, and Not Impactful) between the node and its parent, indicating how impactful the node is to support or oppose its parent claim. We preprocess the original dataset to build one dataset with 3691 argument trees for the impact score of the support relation and one dataset with 3676 argument trees for the rebuttal relation and train models for them separately. To make different classes more balanced, we also incorporate a data resampling mechanism to let each class’ samples occur multiple times according to the ratio of the maximum class count to the number of samples in this class. We then train the base models in the classification task for 3 epochs. Finally, our trained reward model gets 0.67 accuracy in predicting the support impactfulness and 0.72 on the rebuttal impactfulness.

### v2

@@ 4.1 Experiment and Evaluation Setup
We use the state-of-the-art multi-agent framework Agent4Debate [undefal] with different backbone LLMs as our baseline. It employs a collaborative architecture with four specialized agents: searcher, analyzer, writer, and reviewer agents. It shows better performance in the LLM-based Debatrix debate metric [undefr] and is preferred by human experts in pairwise comparison. It uses the stage-specific prompts from the AIDebater 2024 competition  [^2] http://www.fudan-disc.com/sharedtask/AIDebater24 and uses Tavily  [^3] https://tavily.com as a search engine for evidence retrieval.
@@
We use Gemini-2.0-flash [undefae] and DeepSeek-V3 [undefu] as the backbone LLM for Agent4Debate and TreeDebater. For a fair comparison, TreeDebater uses the same Tavily APIs for evidence retrieval before the debate and adopts the same stage-specific prompts, adding the necessary modifications to incorporate the tree information for planning, as shown in Appendix G. To ensure that the debate systems follow the time constraint, we add the rough word budget in the stage-specific prompt for both models. After transferring the statement to the audio via OpenAI TTS, we apply a hard cut on the debate audio: the audio will be trimmed from the last sentence before the time limitation. We train two separate LLaMA-based reward models, specifically with the base model being meta-llama/Llama-3.2-3B-Instruct, using the Kialo dataset [undeff] for $r_{s}$ and $r_{a}$. The Kialo dataset consists of claim texts for 741 controversial motions with three categories (Impactful, Medium Impactful, and Not Impactful). During inference, we use the weighted category as the final prediction to provide a more fine-grained score. The other implementation details are in Appendix B.3.
@@
Stage-level Comparison: Head-to-Head Human Evaluation We design a controlled head-to-head setting for stage-level comparison. We first have a debate competition between the two debate systems with a random stance assignment. Then we use the debate process before a target stage as the context, and let each debate system generate one version for this target stage. We then let the participant compare the two versions based on the debate process. For example, we keep the same Pro’s Opening and ask Agent4Debate and TreeDebater to generate the Con’s Opening. This head-to-head comparison provides a detailed and focused evaluation of each stage by alleviating the noise introduced by other factors. Specifically, the annotator will provide persuasiveness scores for both versions and also mark which side performs better in each stage. We randomly sample 10 (motion, stance assignment) settings for each stage and get the corresponding debate context, , producing 120 stage-level comparisons. Then we let each debate system generate one statement for the target stage. The audio of the debate context and the two versions of the statement are presented to the participants, where we randomly set the order of these two versions to avoid the potential order bias.
@@
Debate-level: End-to-End Human Votes In addition, we follow Oxford-style debating to assess the debater’s capability in a full debate. We ask the participants to vote before and after listening to the full 20-minute debate, and provide the persuasiveness score for each stage. We flip the stance assignment and take the average of two competitions for each motion to avoid the potential prior bias on the motion. We randomly sample 4 motions for Gemini and 7 motions for DeepSeek. We conduct two debates on each motion with the original and swapped sides. For each debate, 3 random participants from our recruitment pool are asked to score each stage and vote before and after the debate. The average score and the opinion shift are used to evaluate the debaters’ performance in this debate. We take the average of the two flipped stance assignments as the debater’s performance on this motion. For example, the opening persuasive score is the average of the opening scores when the debaters acted as the Pro and Con side, respectively. Results are shown in Table 2.
Recruitment and Quality Control We recruit audience members from the online research platform Prolific  [^4] https://www.prolific.com/. The evaluation takes around 30 minutes per stage. The compensation is \$10 for each case after they complete the evaluation. We recruited 212 participants in total for our debate, and all participants are located in the US when they conduct the evaluation. The screenshot of our evaluation platform is shown in Figure 5 and 6, and the demographic features of our participants are shown in Figure 7. For the stage-level head-to-head evaluation, we calculate the annotators’ agreement on the better version. The annotators achieved 60.7% agreement on average, indicating moderate consensus. This agreement indicates a reliable human evaluation result, considering the subjectiveness in debate evaluation.  [^5] In debate-level human evaluation, the winner of the debate is decided based on how many annotators shift towards it, and annotators do not need to agree with each other. Therefore, annotators’ agreement is not applicable here.
@@
Before human evaluation, we perform validity checks to assess the quality of the debate competition generated. Only those that follow the correct debate format are used for human evaluation. The results of the valid checks are shown in Appendix E. In general, TreeDebater can always generate format-valid and time-valid statements. However, only 77% debates are valid in the Gemini version of Agent4Debate. Besides, Agent4Debate always exceeds the time limitation, especially in the closing stage. We put the motion list (Table 5), persuasiveness score rubric (Figure 5), the recruitment details, and a screenshot of our debate evaluation platform (Figure 6) in Appendix C.
@@
Table 1: Scalar persuasiveness score and win rate in head-to-head human evaluation. A higher score indicates the statement is more persuasive. Win:Tie:Lose indicates the number of cases that TreeDebater wins, gets a tie, or loses in the pairwise comparison. The standard deviation results are put in Table 7 in the appendix because of the length limit.
@@
4.2 Experiment Results
TreeDebater has a higher average persuasiveness and win rate in 11/12 stage-level comparison. As shown in Table 1, the advantage is more significant when we shift from Gemini to DeepSeek. We can find a clearer win in the stage-level win rate. When participants are asked to choose the better version, TreeDebater is preferred 1.5x and 2.5x times than the baseline. One reason could be that participants are more likely to choose scores of 3 or 4, leading to a small performance gap in the persuasiveness score.
TreeDebater outperforms the baseline with a 3.5x and 1.3x higher opinion shift win rate in Gemini and DeepSeek. Table 2 indicates that TreeDebater makes the LLM a more persuasive debater that can change the audience’s opinion in the competitive debate. Besides, the debate systems with the DeepSeek backbone have higher persuasiveness scores than Gemini, indicating that DeepSeek has superior capability in argumentation. Moreover, our TreeDebater shows a significant improvement in the persuasiveness score at all stages. We find that this is because the poor performance of the Gemini-based baseline in the early stage affects the participants’ impression of it and makes them more critical in the following stage.
On the other hand, we find that the audience focuses more on the personal belief in the motion rather than the debate strategies if the two sides have reasonably good performance in the debate. For example, the Agent4Debate baseline already shows an average persuasiveness score of 3 or more. In its pairwise comparison with TreeDebater, we find that in the 7 motions we used in our DeepSeek experiments, the Con side always wins in 3 of them and the Pro side always wins in one of them, regardless of the stance assignment. This leads to a comparable persuasiveness score after taking the average of the two flipped assignments, as shown in Table 2. In such a scenario, the head-to-head evaluation provides more insight into the performance difference by focusing more on the debate strategies used under the same debate context.
@@ 4.2 Experiment Results
Gemini Agent4Debate $2.96_{\pm 0.35}$ $2.92_{\pm 0.25}$ $2.98_{\pm 0.29}$ $0.13$
TreeDebater $\textbf{3.60}_{\pm 0.27}$ $\textbf{3.58}_{\pm 0.25}$ $\textbf{3.54}_{\pm 0.29}$ 0.46
DeepSeek Agent4Debate $3.73_{\pm 0.24}$ $3.81_{\pm 0.15}$ $\textbf{3.57}_{\pm 0.23}$ $0.30$
TreeDebater $\textbf{3.87}_{\pm 0.33}$ $71_{\pm 0.28} 3.3$$38_{\pm 0.25}$ 0.40
@@ 4.3 Analysis
Table 3: Ablation Studies on Stage-level Head-to-Head Human Evaluation
Framework Opening Rebuttal Closing
TreeDebater 3.50 3.50 3.75
TreeDebater w/o Rehearsal Tree 3.00 3.25 3.50
TreeDebater w/o Rehearsal & Debater Flow Tree 3.00 3.00 3.50
@@
           [Uncaptioned image] Figure 2: Action distribution in the rebuttal stage. We extract the Debate Flow Tree from human debates and categorize the distribution of actions. The actions are less diverse in the baseline and TreeDebater w/o $T_{d}$.                      [Uncaptioned image] Figure 3: Percentage of actions that can be found in the Rehearsal Trees. The Rehearsal Trees of both sides contribute to the hit rate.
@@
Fine-grained Ablation Studies for Impact on Persuasiveness We further conduct the head-to-head evaluation on the ablation of the Rehearsal Tree and Debate Flow Tree. We select two new motions and use Gemini model as the backbone. The persuasiveness scores are shown Table 3. It shows clear performance degradation when removing components, confirming their necessity for better debate quality.
@@ 5 Conclusion and Discussion
Limitation and Broader Impacts
This paper introduces a TreeDebater to enhance LLM persuasiveness in competitive debates through strategic argumentation and resource management, with broader potential to improve information efficiency in limited-communication settings and strategic planning in games lacking clear winning conditions. However, the study’s conclusions are potentially limited by the scale of human evaluations and the inherent subjectivity of debate performance assessment, despite the authors’ best efforts to mitigate bias through comparative setups; future work could involve larger, more diverse evaluation groups.
@@ Acknowledgments
This work was supported in part by NSF grant IIS-2046640 (CAREER). This project is also supported by the OpenAI gift and the Gemini Academic Program. Xinran Zhao is funded by the ONR Award N000142312840. The authors thank Elizabeth Ma and the Mandarin debate team at HKUST for providing feedback on the debate system and human expert debaters, including Kexun Zhang, Wenting Yang, Yujia Huang, Yue Zhang, Yanran Bao, and Chinmaya Leli, for their valuable feedback.
@@ Appendix A Debate Terminology
We introduce the debate terminologies used in this paper. A motion in the debate is a statement of opinion or proposition to argue. The assigned stance $s$ on the debate motion can be Pro or Con, indicating supporting or opposing the motion, respectively. A claim $c$ is a proposition or assertion about a motion, which should be proved through argumentation. The main claims refer to the most important claims that the debater proposes during the opening stage to support their stance. An argument $x$ is a set of reasons or evidence to support a claim, typically consisting of premises that lead to a conclusion. A statement is the whole passage presented in one debate stage, including a set of claims, arguments, and supporting evidence  [^6] It is also called a speech [undefad] or a discourse [undefak] in some literature.. There are several typical actions that the debater can take in each stage: propose a new claim, reinforce its existing claims, attack the opponent’s claims, and rebut the opponent’s attack. A battlefield refers to a conflict zone in the debate, where two sides contest on a specific point. It includes several rounds of propose attack defense.
@@ B.1 Debate Flow Tree Algorithms
2if actionaction is "propose" then
3   Create node $x^{\prime}$ with claimclaim and arguargu;
@@ 5   Modify $x^{\prime}$ status to "proposed" ;
8Find node $x$ similar to targettarget by $\theta$ ;
9 Increase $x$ visit count by 11 ;
@@
11if actionaction is "reinforce" then
12   Update $x$ with arguargu ;
@@
15if actionaction is "rebut" or "attack" then
16   Create node $x^{\prime}$ with claimclaim and arguargu;
@@ 19   Modify $x^{\prime}$ status to "proposed" ;
Algorithm 2 Update Debate Flow
@@ 5   Initialize retrieval set $R$ ;
7    if actionaction is "propose" or "reinforce" then
8         Find node $x^{l}$ similar to targettarget on the current side by $\theta$;
@@ 9         Add $x^{l}$ and its $k$-step lookahead strength $f_{k}(x^{l})$ to $R$ if exists ;
11    else if actionaction is "attack" or "rebut" then
12         Find node $x^{l}$ similar to targettarget on the opposite side by $\theta$;
@@ 14            Add $x^{l+1}$ and its $k$-step lookahead strength $f_{k}(x^{l+1})$ to $R$ if exists ;
16  Add the tuple actionaction, targettarget, $R$ to $\mathcal{A}_{R}$
@@ B.2 Search Algorithm in Speech Time Control
As mentioned in the main text, our goal is to generate a coherent speech within the time range $t_{l},t_{r}]$ with the chosen actions from the Debate Flow Tree. Note that the number of words is easier to control than the entire speech length for the large language model generation, and the passage word length has a relatively good correlation with the speech length. We can formulate the problem as that we want to call $f(n)$ minimal times to find an output that has speech time in the range $t_{l},t_{r}]$, where $f$ represents an LLM inference where the speech length of $f(n)$ is correlated to $n$.
Note that if at any time the generated speech time ranges in $t_{l},t_{r}]$, we can directly finish the whole process. In our implementation, we first try to establish the boundary word counts for binary search $n_{l},n_{r}]$, where $n_{l}$ produces a speech shorter than $t_{l}$ and $n_{r}$ produces one longer than $t_{r}$. Specifically, we make an initial query that can either become $l$ or $r$, so that we can find the other endpoint by iteratively doubling rlr-l. With this initial interval determined, we can then iteratively refine this interval through binary search, selecting the midpoint and reducing the interval range by half based on the resulting speech duration.
@@ B.3 Reward Model Training Details
We train two separate LLaMA-based reward models, specifically with the base model being meta-llama/Llama-3.2-3B-Instruct, using the Kialo dataset [undeff] for $r_{s}$ and $r_{a}$. The Kialo dataset consists of claim texts for 741 controversial motions. For each motion, it maintains an argument tree. The root of the tree is the motion, and each node is a claim text with its counter-arguments or subsequent arguments as its children. There are human annotations for the impact score (Impactful, Medium Impactful, and Not Impactful) between the node and its parent, indicating how impactful the node is to support or oppose its parent claim. We preprocess the original dataset to build one dataset with 3691 argument trees for the impact score of the support relation and one dataset with 3676 argument trees for the rebuttal relation and train models for them separately. To make different classes more balanced, we also incorporate a data resampling mechanism to let each class’ samples occur multiple times according to the ratio of the maximum class count to the number of samples in this class. We then train the base models in the classification task for 3 epochs. Finally, our trained reward model gets 0.67 accuracy in predicting the support impactfulness and 0.72 on the rebuttal impactfulness. During finetuning, we formulate the original data point with the following prompt and let the llama predict the impact category. We view it as a classification task and use LlamaForSequenceClassification as the backbone architecture and the cross-entropy as the loss function. The instruction format is shown below.
Instruction format of LLaMA-based reward models You are given a chain of arguments, each one supporting or attacking the previous one.
The previous arguments are: [context]
The second last one is: [claim 1]
The last one is: [claim 2]
Now you need to determine the impact of the last one on the second last one, given their relationship [support/attack]. Output only a number among 0, 1, or 2 in your response. 0 means not impactful; 1 means medium impactful; 2 means impactful.
After applying $r_{a}$ and $r_{s}$ to get the strength score by Eqn. 2, the Rehearsal Tree will look like:
One example in Rehearsal Tree Level-0 Root Claim: "claim": "Removing the debt ceiling benefits future generations.", Scores: Support Score: 1.6, Strength: 0.9
Level-1 Opponent’s Attack: "claim": "The debt ceiling, while imperfect, compels fiscal responsibility, safeguarding future generations from unsustainable debt burdens.", Scores: Attack Score: 1.3, Strength: 0.8
Level-2 Your Rebuttal: "claim": "The debt ceiling does not ensure fiscal responsibility; it merely invites fiscal crises.", Scores: Attack Score: 1.4, Support Score: 1.6, Strength: 0.6
Level-3 Opponent’s Attack: "claim": "The debt ceiling’s ’manufactured crises’ are preferable to unchecked spending, which is a greater danger.", Scores: Attack Score: 0.8, Support Score: 0.8, Strength: 0.8
Level-3 Opponent’s Attack: "claim": "Our budgeting process requires improvement, but eliminating the debt ceiling is not the right solution.", Scores: Attack Score: 1.0, Support Score: 0.8, Strength: 0.9
Level-3 Opponent’s Attack: "claim": "The debt ceiling does not need to be a ’political weapon’ and can be reformed to work more effectively, rather than eliminated.", Scores: Attack Score: 1.2, Support Score: 1.0, Strength: 1.1